OpenStudy (anonymous):

(2x)^-2 y^3 over x^3 y^-2 simplify and help me with the steps please

jhonyy9 (jhonyy9):

y3/4x2/x3/y2 = y3/4x2 * y2/x3 =y5/4x5

OpenStudy (anonymous):

perfect man i just god it at the last minute

jhonyy9 (jhonyy9):

thank you